OpenStudy (chrisplusian):

The current entering the positive terminal of a device isi(t)=(6e^{-2t} times 10^{-3}) A and the voltage across the device is v(t)=(10\frac{ di }{ dt })V. Determine the energy absorbed in 3s.

OpenStudy (chrisplusian):

\[i(t)=(6e^{-2t} \times 10^{-3})A\]

OpenStudy (chrisplusian):

\[ v(t)=(10\frac{ di }{ dt })V\]

OpenStudy (vincent-lyon.fr):

Instantaneous power received by the device is \(p(t)=v(t).i(t)\) So energy received is : \(E=\int_{t=0}^{3s}p(t) \,dt \)